2x-5y= -3
2x+5y=4
determine whether each pair represents perpendicular lines


Answer by fcabanski(1391)

The slope-intercept form of an equation is useful for finding the slope, and for comparing lines (parallel or perpendicular).


The equation is y=mx+b where m is the slope. Two lines are perpendicular if their slopes are negative reciprocals of each other. That is, if the product of the slopes is -1.


2x - 5y = -3


Subtract x from both sides.


-5y = -2x - 3


Divide both sides by -5.


y = 2/5x + 3/5.


In order for the second line to be perpendicular to this line, the second line's slope must be -5/2 because 2/5 * -5/2 = -1.


2x + 5y = 4


Subtract 2x from both sides.


5y = -2x + 4


Divide both sides by 5.


y = -2/5x + 4/5


-2/5 * 2/5 = -4/10. That's not -1. The lines are not perpendicular.

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com